OpenStudy (anonymous):

Write the arithmetic sequence that has 3 arithmetic means between -1/4 and -19/12

OpenStudy (anonymous):

I don't know, sorry..

OpenStudy (anonymous):

Nobody seems to be able to answer this question :(

OpenStudy (lalaly):

If the arithmetic sequence has 3 arithmetic means between the 1st and last terms ... there are a total of 5 terms so n=5 a=-1/4 --> 4d-1/4=-19/12 so d=-4/12 and a=-1/4 =-3/12 so the sequence is -3/12 , -7/12, -11/12, -15/12, -19/12

OpenStudy (anonymous):

lana is the best ;) lol <3

OpenStudy (lalaly):

not the best :D thanks sarah ♥ ♥

OpenStudy (anonymous):

haha ur right i am! ;) jkjkj haha <3 <3 ur da besttt lanaaa

OpenStudy (lalaly):

ofcourse you are ♥ ur the sweetest:)

OpenStudy (anonymous):

haha nahh the dorkiestt XD